OpenStudy (anonymous):

what is 1/2*2√3*3√3??

OpenStudy (anonymous):

\[{1 \over 2} \times 2 \sqrt{3} \times 3 \sqrt{3} ?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[{ 1 \over 2 } \times 2 \times \sqrt{3} \times 3 \times \sqrt{3}\] \[{1 \over 2 } \times 2 \times 3 \times \sqrt{3} \times \sqrt{3}\] \[3 \times 3 = 9\]